# David Roodman Open Philanthropy San Francisco, CA

### v3

@@ Abstract
The julia package integrates the Julia programming language into Stata. Users can transfer data between Stata and Julia, issue Julia commands to analyze and plot, and pass results back to Stata. Julia’s econometric ecosystem is not as mature as Stata’s or R’s or Python’s. But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms. For example, the boottest program for wild bootstrap–based inference (Roodman et al., 2019) and fwildclusterboot for R (Fischer and Roodman, 2021) can both call the same Julia back end. And the program reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls a Julia package for tenfold acceleration on hard problems. reghdfejl also supports nonlinear fixed-effect models that cannot otherwise be fit in Stata—though preliminarily, as the Julia package for that purpose is immature.
@@ 1 Introduction
Whenever you issue a command to perform a statistical calculation, you initiate a chain of events of bewildering complexity: your request cascades through layers of code crafted over decades by thousands of programmers in multiple languages, as it is expanded into millions or billions of computational steps—steps that are executed, if you’re lucky, in the time it takes to read this sentence.
In Stata, for example, a call to regress or gsem starts out in the ado programming language. The ado code translates your command into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language. In modern Stata, an ado program can even call Python or Java. The called code will probably invoke the LAPACK library, which is written in Fortran, to invert matrices or solve linear systems.
@@
For creators of computationally intensive software, Julia has many strengths. More so than with Stata (but less so than with Python), if you need a tool to perform a fundamental task such as exploiting GPUs, there’s a good chance someone has written and posted it. Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development and maintenance for speed of execution. (Python and C++ is a common two-language combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that machine code for reuse. A final strength is that, like LISP, Julia can treat code as data. This feature opens the door to meta-programs that can automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
Julia has disadvantages. While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit standard models, and generate plots—many packages are immature or poorly documented. Standards for storing and returning results are less well developed than in Stata. Much less statistical functionality is available than in Stata or R, or even Python.
Since Julia is a great home for implementing and optimizing numerical software, but not (yet) a great home for the end user, a sweet spot for the new language is back end development. Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms. The core work can augmented and optimized in one place, rather than separately on each platform.
The julia package links Stata to Julia. It follows in the footsteps of projects to link statistical software to other languages. StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003. It added Mata in Stata 9. Fiedler (2012) envisioned running Python inside of Stata. Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a C++ plugin for Stata. Perhaps inspired by Fiedler, Stata 16 fully integrated Python. Java plugins became possible in Stata 13 and the Java prompt debuted in the Stata results window in version 17. Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk. Outside of Stata, the JuliaCall (Li, 2019) and JuliaConnectoR (JuliaConnectoR) packages link R to Julia. Another package named JuliaCall does the same for Python. [^1] See juliapy.github.io/PythonCall.jl/stable.
@@
  • The jl: command for Stata, which prefixes single Julia lines or, when typed by itself (with or without the colon), starts a Julia session in Stata.
@@
The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package for Stata.
Two user-written Stata packages call Julia via julia. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata implementation I originally wrote for boottest. And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016). reghdfejl calls the Julia package FixedEffectModels.jl, whose lead developer is Matthieu Gomez. reghdfejl also has the ability, in beta, to fit nonlinear HDFE models, such as the logit and negative binomial, via Johannes Boehm’s GLFixedEffectModels.jl. In the last case, using Julia brings new functionality to Stata. In all cases, calling Julia cuts runtime on hard problems, sometimes by an order of magnitude.
Stata users can thus benefit from Julia and julia without using them directly. Still, this paper explains the new julia package. It starts by demonstrating how to run Julia from Stata, in section 2. Section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia. Section 4 gives tips and warns of some limitations. Section 5 more thoroughly describes the package’s Stata commands and Julia functions.
@@ 2.1 Installation and first examples
Julia is an interpreted language in the sense that you can only run programs written in it if you have installed it. [^2] PackageCompiler.jl can build stand-alone Julia apps, but the files take hundreds of megabytes because they contain complete copies of Julia and its libraries. In contrast, compiled C++ and Fortran programs run on computers without a C++ or Fortran compiler.
The preferred way to get Julia is to download the juliaup installation manager. You do that in Windows by installing Julia (not juliaup) from the Windows Store. [^3] The instructions might change. See github.com/JuliaLang/juliaup#installation for the latest. In Linux and macOS, you can instead start Stata and type
  • ! curl -fsSL https://install.julialang.org | sh -s -- -y
(including the “!” at the beginning). In Linux, you should then restart Stata and the shell to trigger the addition of Julia’s location to the PATH environment variable.
@@
This operation may take several minutes. In Windows, many mysterious console windows may appear and disappear as required libraries are downloaded and pre-compiled. In macOS and Linux, output from the installation process may appear in terminal windows. For on first use, jl installs two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on. DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets, i.e., rectangular grids whose columns have names and potentially different data types. CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.
@@
In the session below, I load Stata’s “auto” data set, extract five variables, run a regression on them, then copy the data to a Julia DataFrame named auto:
@@
jl save is not a Julia command. It is a subcommand of the Stata command jl, which is part of the julia package. If an object named auto already exists in the Julia environment, it is overwritten by the incoming copy of the Stata variables.
@@
Along the top are the column names, and below them the types. Int16 means a 16-bit integer, equivalent to Stata int. Float32 corresponds to Stata float. The question marks at the ends of the type names indicate that the types have been modified to allow the special value missing. By default, jl save made all the columns accept missing.
jl save also converted the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector. While this Julia version of the variable is also stored as a column of integers—1’s and 2’s—it presents as a text variable. [^4] Julia categorical variables are not quite comparable to Stata variables with data labels. Commands involving the Julia variable refer to display values such as “Domestic,” not to numeric codes. The type name “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
@@
  • jl> . jl AddPkg FixedEffectModels  # assure FixedffectModels installed
    jl> . using FixedEffectModels      # load it for use
@@
The Stata and Julia results match. But that Julia regression command needs explaining. First some cosmetic comments:
  • Because the regression command was long, I broke it in two. After I hit the enter key at the end of the first line, jl detected that the line was not grammatical by itself and printed “...” to prompt me to continue.
@@
The command line contains deeper complexity. In Julia, as in R and Python, every regression command is a statement in a full-fledged programming language. This tends to complicate the syntax for even basic operations. It is hard to implement an R or Python line as clean as regress y x. In the Julia command above, term(:price), term(:headroom), …, etc., create instances of the object type Term to represent variables within a statistical model. The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and makes available to the user. The content of each Term is just a symbol, such as :price. What give Term’s life are the operations that StatsModels.jl defines on them. The + operator is repurposed (“overloaded”) to form lists (tuples) of Term’s. The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides, just as in a linear regression model. In the example, an entire FormulaTerm is passed as the second argument to reg(). In this way, the primitives of a flexible programming language are composed to communicate a statistical model.
@@
Yet a powerful feature of Julia makes it practical to ignore most of this complexity. The regression command above can be simplified with a macro. Where macros in Stata are strings, in Julia they are code for manipulating code: meta-programs. StatsModels.jl defines a macro called @formula, which lets one write models more simply. Thus:
@@
    [identical output omitted]
Before this new command is compiled and executed, @formula turns it into something closer to the earlier regression command. [^6] Johannes Boehm’s experimental Douglass.jl package includes macros to translate Stata-like commands into Julia ones. An example: d"bysort :Species : egen :z = sum(:SepalLength) if :SepalWidth .> 3.0".
@@ 2.2 Built for speed
We saw in the previous example that one can start an interactive Julia session inside Stata by typing jl (and end it with exit()). But the interactive mode is not accessible to do files. To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line. It is typographically, not syntactically, split across multiple lines with Stata’s continuation token, ///. When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon, as illustrated above.
@@
Both delegate the computation to built-in matrix operations, which are fast because they are implemented in compiled C or Fortran. However, both contain a subtle inefficiency. They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$, before summing their rows and dispensing with them. Since in our example ${\bf X}$ is $10^{7}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each. Allocating and deallocating memory itself takes time. Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are written to and read from the computer’s main memory. In contrast, the triply-looped Julia code demands essentially no temporary storage.
This inefficiency of Mata and Python can be seen as an artifact of the two-language problem Julia was designed to solve. The best one can do in those languages is rely on a set of canned routines, such as for matrix multiplication, that are written in another language (C++ or Fortran), routines that, while fast on their own terms, are nearly impossible to revise, yet not optimized for the purpose at hand.
@@
On a Windows laptop with an Intel i9-13900H processor, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python. In jl, the data copying is performed by C++ code that is multithreaded across columns.
@@
The Julia function is especially slow on first use because that is when @tturbo performs its magic. On later calls, Julia is fast: it takes 0.05 seconds, which works out to about 40 billion floating point operations per second. Mata takes 0.31 seconds and Python, 0.48. To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above. The Julia package manager would run the script after installing the package in order to trigger immediate compilation. Julia would save the generated machine code for future use.
@@ 3.1 Estimation with high-dimensional fixed effects
In spirit, reghdfejl is a slot-in replacement for reghdfe, with the same syntax, output, and e() return values. In practice, the two differ outwardly as well as inwardly. As reghdfjl depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience. reghdfejl does not support “group fixed effects.” And it does not adjust the reported degrees of freedom for any collinearity of the fixed effects with each other or with the error clustering. So it sometimes reports slightly larger standard errors. While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007). As a result, it can only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic. reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and then transfer control to ivreg2. But in this mode, reghdfejl is no faster than ivreghdfe. For on hard problems, ivreg2 accounts for most of the run time.
@@
specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from groups defined by year. One could attain the same results with bootstrap:reghdfe or bs:reghdfe. But reghdfejl implements vce(bs) in a radically faster way. After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times in a bootstrap sample, and so on.).
To demonstrate reghdfejl’s speed, I create a data set with 10 million rows and two sets of fixed effects. I run the same regression with reghdfe and reghdfejl, as well as with the built-in areg command, which in StataNow 18.5 gained the ability to absorb multiple fixed effects. While reghdfe and reghdfejl can cluster standard errors on multiple variables (CGM_2011), areg cannot, so the example is restricted to one-way clustering:
@@
Using Stata’s bs prefix command with reghdfe takes 69 seconds. Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 12 copies of Stata [^8] See github.com/gvegayon/parallel. cuts the time to 19 seconds. reghdfejl takes just 1.9 seconds. But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 12 worker copies of Julia and triggers some just-in-time compilation in each. The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers reqested by the procs() suboption of the vce() option does not rise further.

### v4

@@ Abstract
The julia package integrates the Julia programming language into Stata. Users can transfer data between Stata and Julia, issue Julia commands to analyze and plot, and pass results back to Stata. Julia’s econometric ecosystem is not as mature as Stata’s or R’s or Python’s. But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms. For example, the boottest program for wild bootstrap–based inference (Roodman et al., 2019) and fwildclusterboot for R (Fischer and Roodman, 2021) can use the same Julia back end. And the program reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects while calling a Julia package for tenfold acceleration on hard problems. reghdfejl also supports nonlinear fixed-effect models that cannot otherwise be fit in Stata—though preliminarily, as the Julia package for that purpose is immature.
@@ 1 Introduction
Whenever you issue a command to perform a statistical calculation, you initiate a chain of events of unfathomable complexity: your request cascades through layers of code crafted over decades by thousands of programmers in multiple languages, as it is expanded into millions or billions of computational steps—steps that are executed, if you’re lucky, in the time it takes to read this sentence.
In Stata, for example, a call to regress or gsem starts out in the ado programming language. The ado code translates your command into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language. In modern Stata, an ado program can even call Python or Java. The called code will probably invoke the Fortran-based LAPACK library to invert matrices or solve linear systems.
@@
For creators of computationally intensive software, Julia has many strengths. More so than with Stata (but less so than with Python), if you need a tool to perform a fundamental task such as exploiting GPUs, there’s a good chance someone has written and posted it. Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development for speed of execution. Python and C++ is a common two-language combination, as is R and C++. Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that machine code for reuse. A final strength is that, like Lisp, Julia can treat code as data. This feature opens the door to meta-programs that can automatically optimize one’s code by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
Julia has disadvantages. While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit common models, and generate plots—many packages are immature or poorly documented. Standards for storing and returning results are less well developed than in Stata. Much less statistical functionality is available than in Stata or R, or even Python.
Since Julia is a great home for numerical software, but not (yet) a great home for the end user, a sweet spot for the language is back end development. Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms. The core can be augmented and optimized in one place, rather than separately on each platform.
The julia package links Stata to Julia. It follows in the footsteps of projects to link statistical software to other languages. StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003. It added the built-in Mata language in Stata 9. Fiedler (2012) envisioned running Python inside of Stata. Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a C++ plugin for Stata. Perhaps inspired by Fiedler, Stata 16 fully integrated Python. Java plugins became possible in Stata 13 and the Java prompt debuted in the Stata results window in version 17. Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk. Outside of Stata, the JuliaCall (Li, 2019) and JuliaConnectoR (Lenz et al., 2022) packages link R to Julia. Another package named JuliaCall does the same for Python. [^1] See juliapy.github.io/PythonCall.jl/stable.
@@
  • The jl command for Stata, which prefixes single Julia lines or, when typed by itself, starts a Julia session in Stata.
@@
The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started with Julia requires nothing more than installing the julia package for Stata.
Two user-written Stata packages call Julia via julia. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata back end I originally wrote for boottest. And the new reghdfejl mimics the reghdfe and ivreghdfe programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016). reghdfejl calls the Julia package FixedEffectModels.jl, whose lead developer is Matthieu Gomez. reghdfejl also has the ability, in beta, to fit nonlinear HDFE models, such as the logit and negative binomial, via Johannes Boehm’s GLFixedEffectModels.jl. In the last case, using Julia brings new functionality to Stata. In all cases, calling Julia cuts runtime on hard problems, sometimes by an order of magnitude.
Stata users can thus benefit from Julia and julia without using them directly. Still, this paper explains the new julia package. It starts by demonstrating how to run Julia from Stata, in section 2. Section 3 introduces applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia. Section 4 gives tips and warns of some limitations. Section 5 more thoroughly describes the package’s Stata commands and Julia functions.
@@ 2.1 Installation and first examples
This operation may take several minutes. In Windows, mysterious console windows may appear and disappear. Julia and the Julia version manager will be installed, if they haven’t been already. The version manager will be configured to guarantee that Julia 1.11, the version currently required by jl, is always available. jl will also install two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on. DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets, meaning rectangular grids whose columns have names and potentially different data types. CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” (Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.)
While Julia’s developers work hard to assure that programs written in older versions work in newer ones too, they do not guarantee backward compatibility in Julia’s low-level, C-based interface, which is what the julia package for Stata links to. Changes to that interface can cause jl to completely crash Stata. (Ask me how I know.) This is why jl is designed to work only with the latest version of Julia at this writing, 1.11. Fortunately, Julia’s version manager makes it easy to maintain several Julia versions at once on a computer. So if you want to use a different version outside of Stata, jl’s need for 1.11 won’t interfere. [^2] If revisions to Julia’s innards break the julia package for Stata, I expect to revise the package to stay up-to-date. The required Julia version may change.
@@
In the session below, I load Stata’s “auto” data set, extract five variables, run a regression on them in Stata, make sure the package I will use for estimation in Julia is installed on my computer, and then copy the data to a Julia DataFrame named auto:
@@
jl AddPkg and jl save are not Julia commands. They are subcommands of the Stata command jl, which is part of the julia package. If an object named auto already exists in the Julia environment, it is overwritten by the incoming copy of the Stata variables.
@@
Along the top are the column names, and below them the types. Int16 means a 16-bit integer, equivalent to Stata int. Float32 corresponds to Stata float. The question marks at the ends of the type names indicate that the types have been modified to allow the special value missing. By default, jl save makes all the columns accept missing.
jl save also converts the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector. The type name “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”. While the Julia version of foreign is also represented internally as a column of integers, it presents as a text variable. All commands involving the Julia variable would refer to display values such as “Domestic,” not to numeric codes as in Stata.
@@
  • jl> . using FixedEffectModels      # load package for use
@@
The Stata and Julia results match. But the Julia regression command needs explaining. First some cosmetic comments:
  • Because the regression command is long, I break it in two. After I hit the enter key at the end of the first line, jl detects that the line is not grammatical by itself and prints “...” to prompt me to continue.
@@
The command line contains deeper complexity. In Julia, as in R and Python, every regression command is a statement in a full-fledged programming language. This tends to complicate the syntax for even basic operations. It is hard to implement an R or Python line as clean as regress y x. In the Julia command above, term(:price), term(:headroom), …, etc., create instances of the object type Term to represent variables within a statistical model. The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and makes available to the user. The content of each Term is just a symbol, such as :price. What give Term’s life are the operations that StatsModels.jl defines on them. The + operator is repurposed (“overloaded”) to form lists (tuples) of Term’s. The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides, just as in a linear regression model. In the example, an entire FormulaTerm is passed as the second argument to reg(). In this way, the primitives of a flexible programming language are composed to articulate a statistical model.
@@
Yet a powerful feature of Julia makes it practical to ignore most of this complexity. The regression command above can be simplified with a macro. Where macros in Stata are strings, in Julia they are code for manipulating code: meta-programs. StatsModels.jl defines a macro called @formula, which lets one write models more simply. Thus, I can write:
@@
      (output omitted )
This is how regressions are usually specified in Julia. Before this new command is compiled and executed, @formula turns it into something closer to the earlier regression command. [^4] Johannes Boehm’s experimental Douglass.jl package includes macros to translate Stata-like commands into Julia ones. An example: d"bysort :Species : egen :z = sum(:SepalLength) if :SepalWidth .> 3.0".
@@ 2.2 Built for speed
We saw in the previous example that one can start an interactive Julia session inside Stata by typing jl (and end it with exit()). But the interactive mode is not accessible in do files. To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line. It is typographically, not syntactically, split across multiple lines with Stata’s continuation token, ///. When putting many Julia commands on what is syntactically one line, most Julia commands not immediately followed by end need to terminate with a semicolon, as illustrated above.
@@
Both delegate the computation to built-in matrix operations, which are fast because they are implemented in compiled C or Fortran. However, both contain a subtle inefficiency. They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$, before summing their rows and dispensing with them. Since in our example ${\bf X}$ is $10^{7}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each. Allocating and deallocating memory itself takes time. Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are transferred to and from the computer’s main memory. In contrast, the triply-looped Julia code demands essentially no temporary storage.
This inefficiency of Mata and Python can be seen as an artifact of the two-language problem Julia was designed to solve. The best one can do in those languages is rely on a set of canned routines, such as for matrix multiplication, that are written in another language (C++ or Fortran), routines that, while fast on their own terms, are not optimized for the purpose at hand and nearly impossible to revise.
@@
We see that on a Windows laptop with an Intel i9-13900H processor, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python. In jl, the data copying is performed by C++ code that is multithreaded across columns.
@@
The Julia function is especially slow on first use because that is when @tturbo performs its magic. On later calls, Julia is fast: it takes 0.05 seconds, which works out to about 40 billion floating point operations per second. Mata takes 0.31 seconds and Python, 0.48. To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above. The Julia package manager would automatically run the script after installing the package in order to trigger immediate compilation. Julia would save the generated machine code for future use.
@@ 3.1 Estimation with high-dimensional fixed effects
In spirit, reghdfejl is a slot-in replacement for reghdfe, with the same syntax, output, and e() return values. In practice, the two differ outwardly as well as inwardly. As reghdfjl depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience. reghdfejl does not support “group fixed effects.” And it does not adjust the reported degrees of freedom for any collinearity of the fixed effects with each other or with the error clustering. So it sometimes reports slightly larger standard errors. While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007). As a result, it can normally only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic. reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and then transfer control to ivreg2. But in this mode, reghdfejl is no faster than ivreghdfe. For on hard problems, ivreg2 accounts for most of the run time.
@@
specifies that 1000 bootstrap replications be performed, split across four CPU threads running in parallel, with the bootstrap data drawn groupwise from groups defined by year. One could attain the same results with bootstrap:reghdfe or bs:reghdfe. But reghdfejl implements vce(bs) in a radically faster way. It copies the primary data just once to Julia. It effects the resampling merely by updating a weight variable, so that, for example, a cluster’s weights are doubled if it is drawn two times in a bootstrap sample.
To demonstrate reghdfejl’s speed on hard problems, I create a data set with 10 million rows and two sets of fixed effects. I run the same regression with reghdfe and reghdfejl, as well as with the built-in areg command, which in StataNow 18.5 gained the ability to absorb multiple fixed effects. While reghdfe and reghdfejl can cluster standard errors on multiple variables (Cameron et al., 2011), areg cannot, so the example is restricted to one-way clustering:
@@
Using Stata’s bs prefix command with reghdfe takes 69 seconds. Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 12 copies of Stata [^6] See github.com/gvegayon/parallel. cuts the time to 19 seconds. reghdfejl takes just 1.9 seconds.